a rectangle has a perimeter of 78 in. its length is 6 in more than twice its length.
Answer by checkley77(12569)

L=2W+6
2L+2W=78
2(2W+6)+2W=78
4W+12+2W=78
6W=78-12
6W=66
W=66/6
W=11 ANS.
L=2*11+6
L=22+6
L=28 ANS.
PROOF:
2*28+2*11=78
56+22=78
78=78